Old Camp Company manufactures awnings for its own line of tents. The company is currently operating at capacity and has received an offer from one of its suppliers to make the 12,000 awnings it needs for $22 each. Old Camp’s costs to make the awning are $11 in direct materials and $6 in direct labor. Variable manufacturing overhead is 80 percent of direct labor. If Old Camp accepts the offer, $41,000 of fixed manufacturing overhead currently being charged to the awnings will have to be absorbed by other product lines.

Complete the incremental analysis for the decision to make or buy the awnings in the table provided below.

Make Buy Net Income:(Increase/Decrease)
Direct Materials
Direct Labor
Variable OverHead
Fixed OverHead
Purchase Price
Total
Answer #1
Make Buy
Direct materials (12,000*11) = 132,000 - 132,000
Direct labor (12,000*6) = 72,000 - 72,000
Voh (72,000*80%) = 57,600 - 57,600
Foh 41,000 41,000 0
Purchase price - (12,000*22) = 264,000 (264,000)
Total 302,600 305,000 (2400)

Comment if you face any issues